The sum of 2 numbers is 45,2/3 of the smaller number is greater than 1/5 of the larger number by 4. Find the two numbers.
Answer by ptaylor(2198)

Let x=the smaller number
Then 45-x=the larger number
Now we are told that:
(2/3)x=(1/5)(45-x)+4 multiply each term by 15
10x=3(45-x)+60 get rid of parens
10x=135-3x+60 add 3x to each side
10x+3x=135-3x+3x+60 collect like terms
13x=195
x=15------------------smaller number
45-x=45-15=30 larger number
CK
(2/3)*15=(1/5)*30+4
10=10
Hope this helps---ptaylor